### v1

# Kalman–Bucy Filtering with Randomized Sensing:

Abstract
Stability analysis of the Kalman filter under randomly lost measurements has been widely studied. We revisit this problem in a general continuous-time framework, where both the measurement matrix and noise covariance evolve as random processes, capturing variability in sensing locations. Within this setting, we derive a closed-form upper bound on the expected estimation covariance for continuous-time Kalman filtering. We then apply this framework to spatiotemporal field estimation, where the field is modeled as a Gaussian process observed by randomly located, noisy sensors. Using clarity, introduced in our earlier work as a rescaled form of the differential entropy of a random variable, we establish a grid-independent lower bound on the spatially averaged expected clarity. This result exposes fundamental performance limits through a composite sensing parameter that jointly captures the effects of the number of sensors, noise level, and measurement frequency. Simulations confirm that the proposed bound is tight for the discrete-time Kalman filter, approaching it as the measurement rate decreases, while avoiding the recursive computations required in the discrete-time formulation. Most importantly, the derived limits provide principled and efficient guidelines for sensor network design problem prior to deployment.
I Introduction
Steady-state analysis of the estimation error reveals how uncertainty evolves and stabilizes over time [1]. In traditional settings, the Kalman filter (KF) provides the optimal linear minimum-variance estimator, with its steady-state covariance governed by a Riccati equation [2, 3, 4]. However, in modern applications such as estimating a spatiotemporal field using a multi-agent sensor network, the sensing process itself is inherently stochastic—sensor locations, measurement timing, and noise may all vary randomly [5, 6, 7]. This randomness makes direct steady-state analysis of the estimation error intractable. Consequently, it is more insightful to adopt a statistical viewpoint, by modeling sensing events as random processes and analyzing their collective impact on estimation performance. Within this stochastic sensing framework, we address a fundamental sensor network design problem constrained by performance requirements: How many sensors, and of what quality, are required to guarantee a prescribed level of estimation accuracy prior to deployment? This motivates two key objectives: (1) to characterize the expected estimation performance analytically, and (2) to establish the fundamental limits of estimation accuracy under stochastic sensing. In this work, we pursue these objectives through both theoretical analysis and application-driven validation.
Prior work has addressed this question in the discrete-time setting. In particular, Gupta [8] analyzed the stochastic sensor selection problem and derived upper bounds on the expected error covariance of the KF under randomly varying measurement availability. While insightful, these bounds are recursive in nature and therefore computationally demanding. In this work, we approximate the measurement process using a continuous-time model rather than directly analyzing the discrete-time recursion. This formulation leads to a differential inequality governing the evolution of the expected estimation covariance, from which the asymptotic behavior of the estimation system naturally follows.
The continuous-time approximation offers two main advantages. First, it admits a closed-form expression for the steady-state covariance, making the analysis computationally much more efficient than the recursive discrete-time formulation. Second, in the context of multi-agent estimation of spatiotemporal Gaussian Process (GP) fields [9], where agents collect noisy measurements at locations drawn from a spatial distribution, the GP structure maps naturally to a linear state-space model, allowing the KF to produce optimal estimates. To quantify estimation uncertainty within this framework, we adopt the clarity, a rescaled form of the differential entropy of a random variable [10]. Leveraging the derived closed-form upper bound and its asymptotic behavior, the proposed approach reveals that the lower bound of the averaged expected clarity is independent of the discretization grid. This property enables the identification of the fundamental limits that govern steady-state estimation performance under stochastic sensing. Consequently, the analysis provides a systematic tool for determining, prior to deployment, (i) the minimal number of agents and (ii) the required sensing capabilities to achieve a prescribed level of estimation accuracy. This directly addresses an optimal sensor resource design problem that cannot be solved efficiently within the discrete-time formulation.
The main contributions are as follows:
  • We develop an explicit closed-form upper bound on the expected estimation covariance for Kalman–Bucy filtering under randomized sensing. Unlike the discrete-time formulation, this bound admits a closed-form steady-state expression, enabling efficient evaluation of estimation performance without recursive numerical computation.
  • We apply the proposed theoretical framework to spatiotemporal field estimation, where the field is modeled as a Gaussian process observed through randomly located, noisy sensors. Within this context, clarity, derived from the differential entropy of a random variable, is employed to characterize estimation uncertainty. Leveraging the established covariance bound, we derive a lower bound on the averaged expected clarity, revealing that steady-state performance is fundamentally constrained by sensing resources and captured by a single composite parameter that unifies the effects of the sensor number, measurement noise, and measurement rate.
  • Numerical simulations confirm that the proposed continuous-time bound becomes increasingly consistent with the discrete-time formulation as the measurement frequency increases. Its closed-form expression provides a practical foundation for sensor number design, offering efficient guidelines for balancing the sensor number, noise level, and sampling rate to achieve a prescribed clarity level.
II Literature Review
A substantial body of work has analyzed the covariance of the KF estimation error, with particular emphasis on its steady-state behavior under various system conditions. For classical time-invariant systems, it is well established that under detectability and stabilizability, the covariance of estimation error converges to a steady state in both discrete- and continuous-time settings [11]. Ni et al. [12] further showed that, for systems with time-varying output matrices, uniform complete observability alone ensures asymptotic stability of the continuous-time KF. However, the systems described above assume deterministic sensing. In contrast, modern sensing environments often deviate from this assumption due to inherent randomness: packet losses, stochastic sensor scheduling, or random sensor placement [7, 8, 13].
Sinopoli et al. [7] pioneered the study of the KF under random measurements by modeling packet losses as a Bernoulli process. They established statistical convergence properties of the expected covariance of estimation error and identified a critical observation-arrival rate, below which the covariance diverges and above which it remains bounded.
Gupta et al. [8] extended this random measurement framework to a more general sensor scheduling problem. They proposed a stochastic sensor selection strategy that activates one sensor at each time step according to a prescribed probability distribution. Using the discrete-time Riccati recursion, they analyzed the evolution of the covariance of estimation error in expectation, deriving upper and lower bounds. Their results further demonstrated, through examples, that these bounds accurately approximate the actual steady-state covariance.
Subsequent research generalized these results to the setting of random measurement matrices. For example, Marelli et al. [5] considered the case where both the measurement matrix and the noise covariance are drawn randomly at each time from a known distribution, and established necessary and sufficient conditions for mean-square stability of the KF. While closely related to our problem, their analysis is limited to stability criteria and does not provide bounds or steady-state characterizations of the error covariance.
Calle and Bopardikar [13] derived probabilistic semidefinite bounds on the covariance of estimation error for randomized sensor selection in Kalman Filtering. Their framework, however, models sensor selection as a one-time random sampling of a subset of sensors that remains fixed thereafter, in contrast to the per-step stochastic scheduling analyzed by Gupta et al. [8].
Early approaches primarily relied on the discrete-time Riccati recursion to track the distribution or expectation of the covariance of estimation error. More recently, efforts have been made to analyze continuous-time process models with discrete-time measurement models [14, 6, 15]. For example, Ahdab et al. [14] modeled measurement arrivals as independent Poisson processes with sensor-specific rates and derived an upper bound on the mean posterior covariance of the continuous–discrete KF along the mean auxiliary state. Similarly, Dasgupta and Tanwani [6] studied continuous-time linear systems with Poisson-sampled observations. They showed that the Kalman–Bucy filter yields piecewise-deterministic covariance trajectories, and derived conditions on the mean observation rate that ensure bounded and convergent expected covariance of estimation error. Their results extend the discrete-time “dropout threshold” in [7]. However, their setting assumes a constant measurement matrix, with randomness arising from the irregular Poisson-distributed observation times rather than from the measurement matrix itself. In contrast, our focus is on a more general problem, akin to the framework of Marelli et al. [5], where the measurement matrix itself is random.
Related work has also addressed continuous measurement models. For example, [16] study sensor scheduling for linear time-invariant systems. The authors derive a convex relaxation of the scheduling problem, yielding a performance lower bound. A lower bound reflects only the best-case performance and offers no guarantee on how the actual performance compares. In contrast, an upper bound constrains the worst-case covariance of the estimation error, providing a concrete performance guarantee. Additionally, [17] consider informative planning of sensing agents over an infinite time horizon for continuous-time linear systems. The covariance bound derived in this work tends to be conservative, as it omits the contribution of the measurement term.
Analysis of the covariance of the estimation error naturally informs the design of the sensing system. Once stability conditions or performance bounds have been established, a practical question arises: how should sensors be deployed or activated to guarantee a prescribed level of estimation accuracy? This leads to sensor scheduling formulations, in which the active subset is selected over time. Prior work has addressed these problems through design optimization grounded in covariance (or its bounds), deriving policies that guide per–time-step sensor activation [18, 14]. Another typical problem is optimal sensor placement: given system parameters, one may wish to determine the minimum number of sensors or the required sensing quality needed to achieve a prescribed uncertainty threshold [19]. Most prior work adopts covariance (or surrogates such as its log-determinant) as the performance metric and designs placement/scheduling policies accordingly. However, covariance measures are often difficult to interpret or compare across different sensing configurations, since they are unit-dependent and different scalarizations (e.g., trace, log-determinant) can rank designs inconsistently. To evaluate the estimation quality, we employ clarity, as introduced in [10, 9], which normalizes the differential entropy of a continuous random variable to the interval $[0,1]$. In this paper, we show that within our closed-form upper-bound framework, the fundamental limits of estimation performance can be explicitly characterized, enabling direct quantification of the sensing resources required to achieve a prescribed uncertainty threshold.
III Problem Formulation
We consider a continuous-time linear system evolving according to a stochastic differential equation (SDE):
$\dot{\bm{x}}(t)=\bm{\bm{A}}\bm{x}(t)+\bm{w}(t),\quad\bm{w}(t)\sim\mathcal{N}(\bm{0},\bm{Q}_{c}),$ (1)
where $\bm{x}(t)\in\mathbb{R}^{n}$ denotes the system state at time $t$, $\bm{\bm{A}}\in\mathbb{R}^{n\times n}$ is the system matrix, $\bm{w}(t)$ is a zero-mean white Gaussian noise with covariance $\bm{Q}_{c}\in\mathcal{S}_{+}^{n}$. The process state is observed by $N_{r}$ sensors whose spatial locations vary randomly over time within a compact spatial domain $D\subset\mathbb{R}^{d}$, where $d$ is the spatial dimension. Specifically, let
$P_{r}(t)=\{\,\bm{r}_{j}(t)\,\}_{j=1}^{N_{r}}$ (2)
denote the set of sensor locations at time $t$, where $\bm{r}_{i}(t)\in D\subset\mathbb{R}^{d}$ denotes the location of the $i$-th sensor. We assume that sensor positions are drawn independently and identically distributed (i.i.d.) from a fixed spatial distribution $\mathcal{D}_{r}$, i.e.,
$P_{r}(t)\;\sim\;\mathcal{D}_{r}.$
Each sensor measures a linear function of the process state corrupted by measurement noise:
$\bm{y}(t)$ $=\bm{H}(t)\bm{x}(t)+\bm{v}(t),\quad\bm{v}(t)\sim\mathcal{N}(0,\bm{V}(t))$ (3)
where $\bm{y}(t)\in\mathbb{R}^{N_{r}}$ is the measurement vector, $\bm{H}(t)\in\mathbb{R}^{N_{r}\times n}$ is the time-varying observation matrix determined by the instantaneous sensor configuration $P_{r}(t)$, and $\bm{v}(t)$ is a zero-mean white Gaussian measurement noise with covariance matrix $\bm{V}(t)$. The measurement noises are assumed independent of the process noise $\bm{w}(t)$. Since all sensors share their measurements, each has access to the same global observation history and maintains a common estimate of the system state, denoted by $\hat{\bm{x}}(t)$.
Given the measurement model above, the optimal estimate of $\bm{x}(t)$ in the mean-square sense is provided by the Kalman–Bucy filter, whose covariance matrix
$\bm{\Sigma}(t):=\mathbb{E}\!\left[(\bm{x}(t)-\hat{\bm{x}}(t))(\bm{x}(t)-\hat{\bm{x}}(t))^{\top}\right]$ (4)
evolves according to the Riccati differential equation
$\dot{\bm{\Sigma}}(t)$ $=\bm{A}\bm{\Sigma}(t)+\bm{\Sigma}(t)\bm{A}^{\top}+\bm{Q}_{c}$ (5)
$-\bm{\Sigma}(t)\bm{H}(t)^{\top}\bm{V}(t)^{-1}\bm{H}(t)\bm{\Sigma}(t),$
where $\bm{\Sigma}(0)=\bm{\Sigma}_{0}\in\mathcal{S}_{+}^{n}$ is the initial covariance matrix.
For notational convenience, we introduce the Riccati operator parameterized by $\bm{G}\in\mathcal{S}_{++}^{n}$ as
$\mathcal{R}_{\bm{G}}(\bm{\Sigma})=\bm{\bm{A}}\bm{\Sigma}+\bm{\Sigma}\bm{\bm{A}}^{\top}+\bm{Q}_{c}-\bm{\Sigma}\bm{G}\bm{\Sigma},$ (6)
where $\bm{\Sigma}\in\mathcal{S}_{++}^{n}$. With this notation, the Kalman–Bucy Riccati equation reads
$\dot{\bm{\Sigma}}(t)=\mathcal{R}_{\bm{G}(t)}(\bm{\Sigma}(t)),\quad\bm{G}(t)=\bm{H}(t)^{\top}\bm{V}(t)^{-1}\bm{H}(t).$ (7)
Because sensor locations $P_{r}(t)$ evolve randomly, both $\bm{H}(t)$ and $\bm{V}(t)$ are random matrix-valued processes, rendering $\bm{\Sigma}(t)$ itself a random matrix process. Our objective is to characterize the expected covariance
$\bar{\bm{\Sigma}}(t):=\mathbb{E}_{P_{r}(t)\sim\mathcal{D}_{r}}[\bm{\Sigma}(t)],$ (8)
and to derive tractable upper bounds on $\bar{\bm{\Sigma}}(t)$ that capture how stochastic sensing influences estimation performance.
For brevity, we will write $\mathbb{E}[\cdot]$ to denote expectation with respect to $P_{r}(t)\sim\mathcal{D}_{r}$.
IV Continuous-Time Upper Bound and Analytical Analysis
IV-A Continuous-Time Upper Bound of Expected Covariance
To establish the continuous-time upper bound, we first derive a set of structural properties of the Riccati operator that enable us to bound $\bar{\bm{\Sigma}}(t)$.
Lemma 1 (Concavity).
The operator $\mathcal{R}_{\bm{G}}$ is concave, that is, for any $\bm{\Sigma}_{1},\bm{\Sigma}_{2}\in\mathcal{S}_{++}^{n}$ and $\lambda\in[0,1]$,
$\mathcal{R}_{\bm{G}}(\lambda\bm{\Sigma}_{1}\!+\!(1-\!\lambda)\bm{\Sigma}_{2})\succeq\lambda\mathcal{R}_{\bm{G}}(\bm{\Sigma}_{1})\!+\!(1-\!\lambda)\mathcal{R}_{\bm{G}}(\bm{\Sigma}_{2})$ (9)
Proof.
See Appendix A-A. ∎
Lemma 2.
For any $\bm{\Sigma}_{1},\bm{\Sigma}_{2}\in\mathcal{S}_{+}^{n}$, let $\bm{K}=\bm{\Sigma}_{1}-\bm{\Sigma}_{2}$. Then,
$\mathcal{R}_{\bm{G}}(\bm{\Sigma}_{1})-\mathcal{R}_{\bm{G}}(\bm{\Sigma}_{2})=\tilde{\bm{A}}\bm{K}+\bm{K}\tilde{\bm{A}}^{\top}$ (10)
where $\tilde{\bm{A}}=\bm{A}-\frac{1}{2}(\bm{\Sigma}_{1}+\bm{\Sigma}_{2})\bm{G}$.
Proof.
See Appendix A-B. ∎
Lemma 3.
Let $\tilde{\bm{A}}:[t_{0},\infty)\to\mathbb{R}^{n\times n}$. Suppose $\bm{K}:[t_{0},\infty)\to\mathbb{R}^{n\times n}$ satisfies
$\dot{\bm{K}}(t)$ $\succeq\tilde{\bm{A}}(t)\bm{K}(t)+\bm{K}(t)\tilde{\bm{A}}(t)^{\top}\quad\forall t\geq t_{0},$ (11)
with initial condition $\bm{K}(t_{0})$. Then,
$\bm{K}(t_{0})\succeq 0\implies\bm{K}(t)\succeq 0\quad\forall t\geq t_{0}.$ (12)
Proof.
See Appendix A-C. ∎
Theorem 1 (Continuous-Time Upper Bound).
Let averaged information matrix
$\bar{\bm{G}}:=\mathbb{E}[\bm{H}(t)^{\top}\bm{V}(t)^{-1}\bm{H}(t)].$ (13)
Define $\bm{\Delta}:[t_{0},\infty)\to\mathcal{S}_{++}^{n}$ as the solution of the Continuous-time algebraic Riccati equation (CARE)
$\dot{\bm{\Delta}}=\bm{A}\bm{\Delta}+\bm{\Delta}\bm{A}^{\top}+\bm{Q}_{c}-\bm{\Delta}\bar{\bm{G}}\bm{\Delta},\quad\bm{\Delta}(t_{0})=\bar{\bm{\Sigma}}(t_{0}).$ (14)
Then, for all $t\geq t_{0}$,
$\bar{\bm{\Sigma}}(t)\preceq\bm{\Delta}(t).$ (15)
Proof.
Given that $P_{r}(t)$ is drawn randomly from $\mathcal{D}_{r}$, we obtain $\bm{G}(t)$ is a continuous matrix-valued random variable. Define $\dot{\bar{\bm{\Sigma}}}=\frac{d}{dt}\mathbb{E}[\bm{\Sigma}(t)]$, therefore,
$\dot{\bar{\bm{\Sigma}}}$ $=\mathbb{E}[\dot{\bm{\Sigma}}(t)]$ (16a)
$=\mathbb{E}[\bm{A}\bm{\Sigma}(t)+\bm{\Sigma}(t)\bm{A}^{\top}+\bm{Q}_{c}-\bm{\Sigma}(t)\bm{G}(t)\bm{\Sigma}(t)]$ (16b)
$=\mathbb{E}[\bm{A}\bm{\Sigma}(t)+\bm{\Sigma}(t)\bm{A}^{\top}+\bm{Q}_{c}-\bm{\Sigma}(t)\bar{\bm{G}}\bm{\Sigma}(t)]$ (16c)
$=\mathbb{E}[\mathcal{R}_{\bar{\bm{G}}}(\bm{\Sigma}(t))]$ (16d)
$\preceq\mathcal{R}_{\bar{\bm{G}}}(\mathbb{E}[\bm{\Sigma}(t)])$ (16e)
where Eq. (16a) holds by linearity of expectation; Eq. (16b) follows by substituting the expression for $\dot{\bm{\Sigma}}$; Eq. (16c) uses the assumed independence of $\bm{\Sigma}(t)$ and $\bm{G}(t)$; Eq. (16d) invokes the definition of the Riccati operator in Eq. (6), and Eq. (16e) follows from Jensen’s inequality together with the concavity of $\mathcal{R}_{\bm{G}}$ established in Lemma 1.
Define $\bm{K}(t)=\bm{\Delta}(t)-\bar{\bm{\Sigma}}(t)$ we have
$\dot{\bm{K}}(t)$ $=\dot{\bm{\Delta}}(t)-\dot{\bar{\bm{\Sigma}}}(t)$ (17)
$\succeq\mathcal{R}_{\bar{\bm{G}}}(\bm{\Delta}(t))-\mathcal{R}_{\bar{\bm{G}}}(\bar{\bm{\Sigma}}(t))$
$=\tilde{\bm{A}}(t)\bm{K}(t)+\bm{K}(t)\tilde{\bm{A}}(t)^{\top}$
where $\tilde{\bm{A}}(t)=\bm{A}-\frac{1}{2}(\bm{\Delta}(t)+\bar{\bm{\Sigma}}(t))\bar{\bm{G}}$, based on the proof in Lemma 2. By Lemma 3, this means that $\bm{K}(t)\succeq 0$ for all $t\geq t_{0}$. ∎
Corollary 1.
Let $\mathcal{L}:\mathcal{S}_{+}^{n}\to\mathbb{R}$ be a linear or concave operator mapping positive semidefinite matrices to scalars. Then
$\mathcal{L}(\bar{\bm{\Sigma}}(t))\leq\mathcal{L}(\bm{\Delta}(t)),\quad\forall t\geq t_{0}.$ (18)
Corollary 2.
Suppose $(\bm{A},\bar{\bm{G}})$ is observable. Then,
$\bar{\bm{\Sigma}}_{\infty}=\lim_{t\to\infty}\bar{\bm{\Sigma}}(t)$ (19)
exists, and
$\bar{\bm{\Sigma}}_{\infty}\preceq\bm{\Delta}_{\infty}.$ (20)
where the $\bm{\Delta}_{\infty}$ is the solution of Eq. (14).
Remark 1.
Theorem 1 provides, to the best of our knowledge, the first continuous-time upper bound on the expected estimation covariance $\bar{\bm{\Sigma}}(t)$ under randomized sensing. For each $t\geq 0$, the covariance is bounded by a deterministic matrix $\bm{\Delta}(t)$, which evolves according to the Riccati differential equation in Eq. (14).
IV-B Closed-form Solution of Upper Bound
To gain further insight into the structure of this bound, we next consider a canonical case in which the system dynamics and process excitation are isotropic, i.e., $\bm{A}=a\bm{I}$ with $a<0$ and $\bm{Q}_{c}=q_{c}\bm{I}$. Under this assumption, the upper-bound Riccati dynamics decouple in the eigenbasis of the averaged information matrix $\bar{\bm{G}}$, reducing to a set of scalar Riccati equations, one along each eigendirection of $\bar{\bm{G}}$ as defined in Eq. (13). This simplification enables a closed-form characterization of the steady-state covariance, as stated in the following theorem.
Theorem 2.
[Closed-Form of Upper Bound] Consider the continuous-time differential equation governing the upper bound $\bm{\Delta}$ as:
$\dot{\bm{\Delta}}\;=\;\bm{A}\bm{\Delta}+\bm{\Delta}\bm{A}^{\top}+\bm{Q}_{c}-\bm{\Delta}\bar{\bm{G}}\bm{\Delta},$ (21)
$\bm{A}=a\bm{I}~(a<0),\;\bm{Q}_{c}=q_{c}\bm{I}~(q_{c}>0),\;\bar{\bm{G}}\in\mathcal{S}_{++}^{n}.$
Then $\bm{\Delta}(t)$ converges to the unique stabilizing solution $\bm{\Delta}_{\infty}$ of the CARE
$0\;=\;a\bm{\Delta}_{\infty}+\bm{\Delta}_{\infty}a+q_{c}\bm{I}-\bm{\Delta}_{\infty}\bar{\bm{G}}\bm{\Delta}_{\infty},$ (22)
given in closed-form by
$\bm{\Delta}_{\infty}=-q_{c}\left(a\bm{I}-\sqrt{a^{2}\bm{I}+q_{c}\bar{\bm{G}}}\right)^{-1}.$ (23)
Equivalently, if $\{\lambda_{i}\}_{i=1}^{n}$ are the eigenvalues of $\bar{\bm{G}}$, then the eigenvalues of $\bm{\Delta}_{\infty}$ are
$\gamma_{i}=\frac{-q_{c}}{a-\sqrt{a^{2}+q_{c}\lambda_{i}}},\qquad i=1,\dots,n.$ (24)
Proof.
Since $\bar{\bm{G}}\in\mathbb{S}_{++}^{n}$, it admits the eigen-decomposition $\bar{\bm{G}}=\bm{U}\Lambda\bm{U}^{\top}$, where $\bm{U}$ is orthogonal and $\Lambda$ is diagonal. Under the isotropic assumption $\bm{A}=a\bm{I}$ and $\bm{Q}_{c}=q_{c}\bm{I}$, Eq. (22) implies that $\bar{\bm{G}}$ and $\bm{\Delta}_{\infty}$ commute, i.e., $\bar{\bm{G}}\bm{\Delta}_{\infty}=\bm{\Delta}_{\infty}\bar{\bm{G}}$. By Theorem 4.5.15 (a) in [20], any two commuting symmetric matrices can be simultaneously diagonalized by the same orthogonal matrix. Hence, $\bm{\Delta}_{\infty}=\bm{U}\bm{\Gamma}\bm{U}^{\top}$, where $\bm{\Gamma}$ is diagonal.
Applying the same factorization to $\bm{\Delta}_{\infty}$, we have $\bm{\Delta}_{\infty}=\bm{U}\bm{\Gamma}\bm{U}^{\top}$, where $\bm{\Gamma}$ is diagonal. Therefore, the equation becomes
$0$ $=\bm{A}\bm{\Delta}_{\infty}+\bm{\Delta}_{\infty}\bm{A}^{\top}+\bm{Q}_{c}-\bm{\Delta}_{\infty}\bar{\bm{G}}\bm{\Delta}_{\infty}$
$=2a(\bm{U}\bm{\Gamma}\bm{U}^{\top})+q_{c}\bm{I}-(\bm{U}\bm{\Gamma}\bm{U}^{\top})(\bm{U}\Lambda\bm{U}^{\top})(\bm{U}\bm{\Gamma}\bm{U}^{\top})$
$=q_{c}\bm{I}+\bm{U}(2a\bm{\Gamma}-\bm{\Gamma}\Lambda\bm{\Gamma})\bm{U}^{\top}$
$=\bm{U}(q_{c}\bm{I}+2a\bm{\Gamma}-\bm{\Gamma}\Lambda\bm{\Gamma})\bm{U}^{\top}$
where the term inside the parentheses is a diagonal matrix. For the equation to hold, each diagonal entry must be zero. Therefore, let $\gamma_{i}=\bm{\Gamma}_{ii}$, $\lambda_{i}=\Lambda_{ii}$, and we have
$q_{c}+2a\gamma_{i}-\lambda_{i}\gamma_{i}^{2}=0,$ (25)
which has the solutions
$\gamma_{i}=\frac{a\pm\sqrt{a^{2}+q_{c}\lambda_{i}}}{\lambda_{i}}.$ (26)
Since we seek the stabilizing solution, we keep only the positive solutions, i.e.,
$\gamma_{i}=\frac{a+\sqrt{a^{2}+q_{c}\lambda_{i}}}{\lambda_{i}}=\frac{-q_{c}}{a-\sqrt{a^{2}+q_{c}\lambda_{i}}}\quad i=1,\dots,n.$ (27)
Thus we have derived the eigenvalues of $\bm{\Delta}_{\infty}$. Using these eigenvalues, we have
$\bm{\Delta}_{\infty}=\bm{U}(\operatorname{diag}(\gamma_{1},...,\gamma_{n}))\bm{U}^{\top}$ (28)
which can be expressed succinctly as
$\bm{\Delta}_{\infty}=\bar{\bm{G}}^{-1/2}\left(a\bm{I}+(a^{2}\bm{I}+q_{c}\bar{\bm{G}})^{1/2}\right)\bar{\bm{G}}^{-1/2}$ (29)
Since $\bar{\bm{G}}\in\mathcal{S}_{++}^{n}$, $(a^{2}\bm{I}+q_{c}\bar{\bm{G}})^{1/2}$ is a closed-form matrix function of $\bar{\bm{G}}$ and thus commutes with $\bar{\bm{G}}^{\pm 1/2}$. Hence,
$\bm{\Delta}_{\infty}$ $=a\,\bar{\bm{G}}^{-1/2}\bm{I}\bar{\bm{G}}^{-1/2}+(a^{2}\bm{I}+q_{c}\bar{\bm{G}})^{1/2}\bar{\bm{G}}^{-1/2}\bar{\bm{G}}^{-1/2}$ (30)
$=\left(a\bm{I}+(a^{2}\bm{I}+q_{c}\bar{\bm{G}})^{1/2}\right)\bar{\bm{G}}^{-1}$
$=-q_{c}\left(a\bm{I}-\sqrt{a^{2}\bm{I}+q_{c}\bar{\bm{G}}}\right)^{-1}.$
Remark 2.
Theorem 2 provides an explicit closed-form expression for the steady-state upper bound on the expected covariance of the estimation error. The solution reveals how the eigenvalues of the averaged information matrix $\bar{\bm{G}}$ directly determine the achievable estimation accuracy, offering a tractable and interpretable characterization of continuous-time estimation performance.
Based on above analysis, understanding what factors determine $\bar{\bm{G}}$ is key to explaining how sensing resources influence estimation performance. This connection motivates the following section, in which we derive a closed-form characterization of $\bar{\bm{G}}$ under a spatiotemporal sensing environment.
V Application to Spatiotemporal Estimation
The continuous-time upper-bound framework developed in the previous section provides a general characterization of estimation uncertainty under stochastic sensing. We now consider the problem of spatiotemporal field estimation, where the process of interest is modeled as a GP evolving continuously in both space and time. To estimate the spatiotemporal field from noisy measurements collected at randomized sensor locations, we employ the  spatiotemporal Gaussian process Kalman filter (STGPKF) proposed in [9]. In this section, we demonstrate that the continuous-time upper-bound results derived earlier can be directly applied to the STGPKF, enabling a principled analysis of the fundamental limits of spatiotemporal estimation performance.
An overview of main notations used in this section is provided in Table I.
TABLE I: Summary of covariance- and clarity-related symbols.
Symbol Definition
Covariance-related symbols
$\bm{\Sigma}$ Covariance of the latent-state estimation error $\bm{s}-\hat{\bm{s}}$.
$\bar{\bm{\Sigma}}=\mathbb{E}[\bm{\Sigma}]$ Expected covariance of the latent-state estimation error under stochastic sensing.
$\bm{\Delta}$ Upper bound of $\bar{\bm{\Sigma}}$.
$\bm{\Pi}$ Covariance of the field-state estimation error $\bm{f}-\hat{\bm{f}}$.
$\bar{\bm{\Pi}}=\mathbb{E}[\bm{\Pi}]$ Expected covariance of the field-state estimation error under stochastic sensing.
$\bm{\Delta}^{\Pi}$ Upper bound of $\bar{\bm{\Pi}}$.
$\bar{\bm{\Sigma}}_{\infty}$, $\bm{\Delta}_{\infty}$ Steady-state $\bar{\bm{\Sigma}}$, $\bm{\Delta}$.
Clarity-related symbols
$q_{i}$ Clarity at the $i$-th spatial location.
$\bar{q}$ Spatially averaged clarity for a given sensor realization.
$\bar{q}_{\mathbb{E}[\Pi]}$ Averaged expected clarity defined in Eq. (60).
$\bar{q}_{\Delta^{\Pi}}$ Lower bound on averaged expected clarity in defined Eq. (61).
$\bar{q}_{\Delta^{\bm{\Pi}}_{\infty}}$ Steady-state lower bound of averaged expected clarity defined in Eq. (64).
V-A Spatiotemporal Environment
We consider a spatiotemporal field $f:\mathbb{R}\times D\to\mathbb{R}$, modeled as a GP with a separable kernel
$f(t,p)\sim\mathcal{GP}\big(0,\,k_{T}(t,t^{\prime})\,k_{S}(\bm{p},\bm{p}^{\prime})\big),\,\bm{p}\in D\subset\mathbb{R}^{d},$ (31)
where $k_{T}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ and $k_{S}:D\times D\to\mathbb{R}$ are continuous, symmetric, positive definite kernels capturing temporal and spatial correlations, respectively. For example, we may use the Matérn-$\tfrac{1}{2}$ kernel in both time and space:
$k_{T}(t,t^{\prime})$ $=\sigma_{t}^{2}\exp\!\left(-\tfrac{|t-t^{\prime}|}{l_{t}}\right),$ (32)
$k_{S}(\bm{p},\bm{p}^{\prime})$ $=\sigma_{s}^{2}\exp\!\left(-\tfrac{\|\bm{p}-\bm{p}^{\prime}\|}{l_{s}}\right),$ (33)
with variances $\sigma_{t}^{2},\sigma_{s}^{2}$ and correlation lengths $l_{t},l_{s}$.
Agents act as sensors that probe this GP field. At a fixed sampling period $\Delta t>0$, measurements are collected by $N_{r}$ agents whose sampling locations vary randomly. Specifically, at the $k$-th timestep ($t_{k}=k\Delta t$), the agent locations are drawn i.i.d. from a spatial distribution $\mathcal{D}_{r}$:
$P_{r}(t_{k})\;\sim\;\mathcal{D}_{r}.$ (34)
At each location, the $j$-th agent acquires a noisy measurement
$y_{j}(t_{k})\;=\;f(t_{k},\bm{r}_{j}(t_{k}))+\eta_{j},\quad\eta_{j}\sim\mathcal{N}(0,\sigma_{m}^{2}),$ (35)
where $y_{j}(t_{k})\in\mathbb{R}$ is a scalar measurement output by agent $j$ at time $t_{k}$, and $\eta_{j}\in\mathbb{R}$ is zero-mean Gaussian noise with variance $\sigma_{m}^{2}$.
V-B Spatiotemporal Gaussian Process Kalman Filter
To estimate the spatiotemporal field $f(t,\bm{p})$ from the randomized sensor measurements defined in Eq. (35), we begin by introducing a finite-dimensional approximation of the spatial domain. For grid spacing $\delta>0$, define the uniform grid
$P_{g}^{\delta}=\{\bm{p}^{\delta}_{1},\dots,\bm{p}^{\delta}_{N_{g}^{\delta}}\}\subset D,$ (36)
where $\bm{p}^{\delta}_{i}\in D\subset\mathbb{R}^{d}$ denotes the location of $i$-th grid point. The total number of grid points is given by $N_{g}^{\delta}:=\frac{|D|}{\delta^{d}}$, where $|D|>0$ is the Lebesgue measure of $D$.
Following previous work in [9], the temporal kernel $k_{T}$ admits an equivalent finite-dimensional SDE representation. At each $\bm{p}^{\delta}_{i}$, let $\bm{s}_{i}(t)\in\mathbb{R}^{n_{k}}$ denote the latent temporal state induced by the temporal kernel, where $n_{k}$ is the state dimension. Stacking yields $\bm{s}(t):=[\bm{s}_{1}(t)^{\top},\ldots,\bm{s}_{N_{g}^{\delta}}(t)^{\top}]^{\top}\in\mathbb{R}^{N_{g}^{\delta}n_{k}}$. The latent state evolves according to
$d\bm{s}(t)$ $=\big(\bm{I}_{N_{g}^{\delta}}\otimes\bm{A}_{0}\big)\,\bm{s}(t)\,dt+\big(\bm{I}_{N_{g}^{\delta}}\otimes\bm{B}_{0}\big)\,d\bm{W}(t)$ (37)
$=\bm{A}\,\bm{s}(t)\,dt+\bm{B}\,d\bm{W}(t),$
where $\bm{W}(t)=[W_{1}(t),\ldots,W_{N_{g}^{\delta}}(t)]^{\top}$ collects independent standard Wiener processes, $\bm{A}_{0}\in\mathbb{R}^{n_{k}\times n_{k}}$ and $\bm{B}_{0}\in\mathbb{R}^{n_{k}\times 1}$ are the state and diffusion matrices for a single grid point determined by the chosen temporal kernel, and $\bm{A}:=\bm{I}_{N_{g}^{\delta}}\otimes\bm{A}_{0}$, $\bm{B}:=\bm{I}_{N_{g}^{\delta}}\otimes\bm{B}_{0}$. The continuous-time process model in Eq. (37) is equivalent to Eq. (1) with the process noise covariance defined as $\bm{Q}_{c}=\bm{B}\bm{B}^{\top}$ according to [21]. For the Matérn-$\tfrac{1}{2}$ kernel ($n_{k}=1$), $\bm{A}_{0}$ and $\bm{B}_{0}$ reduce to scalars, corresponding exactly to the isotropic case discussed earlier, so the result of Theorem 1 applies directly.
The spatial correlations among grid points are characterized by the kernel matrix
$\bm{K}_{gg}^{\delta}\in\mathbb{R}^{N_{g}^{\delta}\times N_{g}^{\delta}},\,\big[\bm{K}_{gg}^{\delta}\big]_{ij}=k_{S}\big(\bm{p}^{\delta}_{i},\bm{p}^{\delta}_{j}\big).$ (38)
For brevity, when the grid spacing $\delta$ is fixed we drop the superscript and write $\bm{K}_{gg}:=\bm{K}_{gg}^{\delta},\,N_{g}:=N_{g}^{\delta},\,\bm{p}_{i}:=\bm{p}_{i}^{\delta}$.
Then, the field state at the grid points are obtained from the latent state via
$\bm{f}(t)=\sqrt{\bm{K}_{gg}}(\bm{I}_{N_{g}}\otimes\bm{C}_{0})\bm{s}=\sqrt{\bm{K}_{gg}}C\bm{s}$ (39)
where $\bm{C}_{0}\in\mathbb{R}^{1\times n_{k}}$ is the local observation matrix that maps the latent temporal state $\bm{s}_{i}(t)\in\mathbb{R}^{n_{k}}$ at grid point $i$ to the scalar field value $f_{i}(t)\in\mathbb{R}$, and $\bm{C}:=\bm{I}_{N_{g}}\otimes\bm{C}_{0}$.
While the process model captures GP dynamics on a discretized grid, agents can acquire noisy measurements at random off-grid locations. To relate the latent grid-based state $\bm{s}(t)$ to these physical observations, we introduce the following spatial kernel blocks:
$\bm{K}_{gr}(\bm{r}(t_{k}))$ $=\big[k_{S}(\bm{p}_{i},\,\bm{r}_{j}(t_{k}))\big]_{i,j},$ (40a)
$\bm{K}_{rg}(\bm{r}(t_{k}))$ $=\bm{K}_{gr}^{\top}(\bm{r}(t_{k})),$ (40b)
$\bm{K}_{rr}(\bm{r}(t_{k}))$ $=\big[k_{S}(\bm{r}_{i}(t_{k}),\,\bm{r}_{j}(t_{k}))\big]_{i,j}.$ (40c)
Conditioning the joint Gaussian over grid and agent locations yields the discrete-time linear measurement model
$\bm{y}(t_{k})=\bm{H}(\bm{r}(t_{k}))\,\bm{s}(t_{k})+\bm{v}(t_{k}),$ (41)
$\bm{v}(t_{k})\sim\mathcal{N}\!\big(\bm{0},\bm{V}(\bm{r}(t_{k}))\big),$
with
$\bm{H}(\bm{r}(t_{k}))$ $=\bm{K}_{rg}(\bm{r}(t_{k}))\,\bm{K}_{gg}^{-1}\,\sqrt{\bm{K}_{gg}}\bm{C},$ (42)
$\bm{V}(\bm{r}(t_{k}))$ $=\sigma_{m}^{2}\bm{I}_{N_{r}}+\bm{K}_{rr}(\bm{r}(t_{k}))$
$-\bm{K}_{rg}(\bm{r}(t_{k}))\bm{K}_{gg}^{-1}\bm{K}_{gr}(\bm{r}(t_{k})).$
Here $\bm{H}(\bm{r}(t_{k}))$ is the sensing matrix and $\bm{V}(\bm{r}(t_{k}))$ is the measurement noise covariance, both determined by the instantaneous sensor locations $\bm{r}(t_{k})$ and the measurement noise variance $\sigma_{m}^{2}$.
For compactness, we adopt the shorthand $\bm{r}_{k}:=\bm{r}(t_{k}),\bm{H}_{k}:=\bm{H}(\bm{r}_{k}),\bm{V}_{k}:=\bm{V}(\bm{r}_{k})$. More generally, for any time-varying quantity $X$ in discrete-time setting, we use $X_{k}:=X(t_{k})$.
For the continuous-time process model and discrete measurement model defined in (37) and (LABEL:eq:measurement_model), we follow the approach in [9], which discretizes the continuous-time dynamics into an equivalent discrete-time state–space representation (see, e.g., [22]):
$\bm{s}_{k+1}=\Phi\,\bm{s}_{k}+\bm{w}_{k},\,\bm{w}_{k}\sim\mathcal{N}(\bm{0},\bm{Q}_{d}),$ (43)
where the transition and process-noise matrices are defined as $\bm{\Phi}:=\bm{I}_{N_{g}^{\delta}}\otimes\Phi_{0},\,\bm{Q}_{d}:=\bm{I}_{N_{g}^{\delta}}\otimes\bm{Q}_{d_{0}}$ with
$\Phi_{0}=e^{\bm{A}_{0}\Delta t},\quad\bm{Q}_{d_{0}}=\int_{0}^{\Delta t}e^{\bm{A}_{0}\tau}\bm{B}_{0}\bm{B}_{0}^{\top}e^{\bm{A}_{0}^{\top}\tau}d\tau$ (44)
As shown in [9], the resulting discrete-time KF provides the optimal state estimate, and Eq. (43)–(LABEL:eq:measurement_model) together define a standard linear discrete-time KF. Let $(\hat{\bm{s}}_{k|k},\,\bm{\Sigma}_{k|k})$ denote the posterior mean and covariance of the latent state at step $k$. Conditioned on the time-varying sensing pair $(\bm{H}_{k},\bm{V}_{k})$, the KF update is:
Update step:
$\bm{S}_{k}$ $=\bm{H}_{k}\,\bm{\Sigma}_{k|k-1}\,\bm{H}_{k}^{\top}+\bm{V}_{k},$ (45)
$\bm{K}_{k}$ $=\bm{\Sigma}_{k|k-1}\,\bm{H}_{k}^{\top}\,\bm{S}_{k}^{-1},$
$\hat{\bm{s}}_{k|k}$ $=\hat{\bm{s}}_{k|k-1}+\bm{K}_{k}\big(\bm{y}_{k}-\bm{H}_{k}\hat{\bm{s}}_{k|k-1}\big),$
$\bm{\Sigma}_{k|k}$ $=\big(\bm{I}-\bm{K}_{k}\bm{H}_{k}\big)\,\bm{\Sigma}_{k|k-1}.$
Prediction step:
$\hat{\bm{s}}_{k+1|k}$ $=\bm{\Phi}\,\hat{\bm{s}}_{k|k},$ (46)
$\bm{\Sigma}_{k+1|k}$ $=\bm{\Phi}\,\bm{\Sigma}_{k|k}\,\bm{\Phi}^{\top}+\bm{Q}_{d}.$
Finally, the estimated posterior field on the grid is
$\hat{\bm{f}}_{k|k}$ $=\sqrt{\bm{K}_{gg}}\bm{C}\hat{\bm{s}}_{k|k},$ (47a)
$\bm{\Pi}_{k|k}$ $=\sqrt{\bm{K}_{gg}}\bm{C}\bm{\Sigma}_{k|k}\bm{C}^{\top}\sqrt{\bm{K}_{gg}}.$ (47b)
For brevity, we adopt the shorthand $\bm{\Sigma}_{k}:=\bm{\Sigma}_{k|k-1}$, $\bm{\Pi}_{k}:=\bm{\Pi}_{k|k-1}$, and $\hat{\bm{f}}_{k}:=\hat{\bm{f}}_{k|k}$. For any symmetric matrix $X\in\mathbb{R}^{n\times n}$, define the linear operator
$\mathcal{T}(X):=\bm{K}_{gg}^{1/2}\bm{C}X\bm{C}^{\top}\bm{K}_{gg}^{1/2}.$ (48)
Applying the KF recursion in Eq. (45) and Eq. (46) to this model yields the STGPKF.
V-C Estimation Performance
To evaluate the estimation performance of the STGPKF, we adopt clarity [9], which rescales the differential entropy of a continuous random variable onto the range $[0,1]$, with higher values indicating greater certainty. The differential entropy $h[\bm{Y}]$ of a continuous random variable $\bm{Y}\in\mathbb{R}^{n}$ is defined as:
$h[\bm{Y}]=-\int\rho(x)\,\log\rho(x)\,dx.$ (49)
The clarity of $\bm{Y}$, denoted $q[\bm{Y}]$, is then defined as:
$q[\bm{Y}]=\left(1+\frac{e^{2h[\bm{Y}]}}{(2\pi e)^{n}}\right)^{-1}.$ (50)
Higher clarity values correspond to lower uncertainty, while lower values indicate higher uncertainty. For a scalar Gaussian random variable $Y\sim\mathcal{N}(\mu,\sigma^{2})$, the clarity simplifies to
$q[Y]=\frac{1}{1+\sigma^{2}}.$ (51)
For a Gaussian vector $\bm{Y}\sim\mathcal{N}(0,\bm{\Pi})$ with covariance matrix $\Pi$, we define the marginal clarity for each component using its variance:
$q_{i}=\frac{1}{1+\bm{\Pi}_{ii}},$ (52)
where $\bm{\Pi}_{ii}$ denotes the variance at location $i$.
V-D Numerical Evaluation of STGPKF and Clarity
TABLE II: Simulation parameters.
Parameter Symbol / Value Unit
Temporal kernel std. $\sigma_{t}=2.0$ km/min
Temporal length scale $l_{t}=60$ min
Spatial kernel std. $\sigma_{s}=1.0$ km/min
Spatial length scale $l_{s}=2.0$ km
Spatial resolution $\delta=0.5$ km
Spatial domain $0,5]\times[0,5]$ km2
Temporal resolution $\Delta t=0.05$ min
Simulation horizon $t_{\max}=2\times 60$ min
Measurement noise std. $\sigma_{m}=2.0$ km/min
Number of agents $N_{r}=1$
Figure 1: Wind field reconstruction and clarity assessment using STGPKF for a single realization with $\bm{\Delta}s=0.25$ and $N_{r}=20$. The red–blue color map denotes wind flow direction, and the color intensity represents the wind speed magnitude.
Figure 2: Spatially averaged clarity $\bar{q}_{k}$ over time for different numbers of agents.
In this section, we validate the STGPKF and assess its estimation performance using numerical simulations. The spatiotemporal field, modeled as a GP with temporal and spatial Matérn-$\tfrac{1}{2}$ correlations, admits the following state-space representation [9]:
$\bm{A}_{0}=\begin{bmatrix}-\frac{1}{l_{t}}\end{bmatrix},\,\bm{B}_{0}=\begin{bmatrix}1\end{bmatrix},\,\bm{C}_{0}=\begin{bmatrix}\sigma_{t}\sqrt{\frac{2}{l_{t}}}\end{bmatrix}$
Unless otherwise stated, the parameters used in all subsequent simulations are listed in Table II.
Figure 1 illustrates a representative realization of sensor deployment, where the agent is randomly placed at each time step. The ground-truth field represents the underlying wind flow, where the red–blue color map encodes the wind velocity and the color intensity corresponds to the wind speed magnitude. The STGPKF framework effectively reconstructs the dynamic wind field, accurately capturing both the direction and magnitude of the true flow. The clarity map further verifies the estimation reliability: higher clarity values indicate lower uncertainty and greater confidence in the reconstructed field.
To evaluate the information content of the entire field, rather than at individual points, we consider the spatially averaged clarity. For a given sensor realization, the spatially averaged clarity at time $t_{k}$ is defined as
$\bar{q}_{k}:=\frac{1}{N_{g}}\sum_{i=1}^{N_{g}}\frac{1}{1+\bm{\Pi}_{ii,k}}.$ (53)
where $\bm{\Pi}_{k}=\mathcal{T}(\bm{\Sigma}_{k})$ denotes the covariance of field-state estimation error.
Figure 2 reports the spatially averaged clarity $\bar{q}_{k}$ defined in Eq. (53) evolves over time. As the number of agents increases, averaged clarity improves monotonically. For example, simulations indicate that deploying 20 agents is sufficient to reach the target clarity level of 0.7 on steady-state. In the following section, instead of relying on simulations to determine the attainable clarity for a given number of sensors, we replace such empirical evaluation with a closed-form upper bound, which allows direct computation of the minimum number of sensors required to meet a prescribed clarity threshold.
V-E Steady-State Performance Analysis for STGPKF
A central design question in spatiotemporal field estimation is- Given the environmental parameters, how many sensors, and with what sensing capabilities, are required to guarantee a desired estimation accuracy prior to deployment? To address this question, our first objective is to analyze the expected estimation covariance under randomized sensing. Specifically, this section develops a steady-state bound for the spatially averaged clarity, providing a tractable basis for analyzing fundamental performance limits.
V-E1 Discrete-Time Upper Bound for STGPKF
We first extend the discrete-time stochastic sensor selection framework of [8], which derives upper and lower bounds on the expected estimation covariance using the Riccati recursion and demonstrates their tightness in steady state. Although their formulation considers low-dimensional systems with single-sensor activation, its principles remain applicable. Here, we generalize this framework to our scenario, STGPKF, enabling analysis under multi-agent randomized sensing.
For notational convenience, define the discrete-time Riccati operator parameterized by $(\bm{H},\bm{V})$:
$\mathcal{R}_{\bm{H},\bm{V}}(\bm{\Sigma})$ $=\bm{\Phi}\bm{\Sigma}\bm{\Phi}^{\top}+\bm{Q}_{d}$ (54)
$-\bm{\Phi}\bm{\Sigma}\bm{H}^{\top}(\bm{H}\bm{\Sigma}\bm{H}^{\top}+\bm{V})^{-1}\bm{H}\bm{\Sigma}\bm{\Phi}^{\top},$
where $\bm{\Sigma}\in\mathcal{S}_{++}^{n}$. With the shorthand $\bm{\Sigma}_{k}:=\bm{\Sigma}_{k|k-1}$, the Kalman covariance recursion in Eq. (45)–(46) can be written compactly as
$\bm{\Sigma}_{k+1}\;=\;\mathcal{R}_{\bm{H}_{k},\bm{V}_{k}}(\bm{\Sigma}_{k}).$ (55)
Note that the covariance $\bm{\Sigma}_{k+1}$ is random at each time step $k+1$, since the sensor deployment is drawn from a distribution, i.e., $P_{r,k+1}\sim\mathcal{D}_{r}$. We aim to evaluate:
$\bar{\bm{\Sigma}}_{k+1}:=\mathbb{E}[\bm{\Sigma}_{k+1}]=\mathbb{E}[\mathcal{R}_{\bm{H}_{k},\bm{V}_{k}}(\bm{\Sigma}_{k})].$ (56)
For each time step $k$, this expectation can be approximated via Monte Carlo sampling using $N$ independent sensor deployments $\{(\bm{H}^{(i)}_{k},\bm{V}^{(i)}_{k})\}_{i=1}^{N}$. If deployment $i$ occurs at step $k$ with probability $\pi_{i}^{k}$, then
$\mathbb{E}[\bm{\Sigma}_{k+1}]\approx\sum_{i}^{N}\pi^{k}_{i}\bm{\Sigma}^{(i)}_{k+1}$. (57)
However, explicitly evaluating this expectation is intractable, so we seek an upper bound. The following recursion, adapted from Gupta et al. [8], provides such a result.
Proposition 1.
[8, Thm. 3] [Discrete-time Upper Bound] Let $N$ be the number of sensor deployments. If the $i$-th deployment is chosen at time step $k$ with probability $\pi^{k}_{i}$ independently, then the expected covariance is upper bounded by $\bm{\Delta}_{k+1}$, where $\bm{\Delta}_{k}$ is given by the recursion:
$\bm{\Delta}_{k+1}$ $=\sum_{i=1}^{N}\pi_{i}^{k}\mathcal{R}_{\bm{H}^{(i)}_{k},\bm{V}^{(i)}_{k}}(\bm{\Delta}_{k})$ (58)
$=\bm{Q}_{d}+\bm{\Phi}\bm{\Delta}_{k}\bm{\Phi}^{\top}$-
$\sum_{i=1}^{N}\pi^{k}_{i}\left[\bm{\Phi}\bm{\Delta}_{k}\bm{H}_{i,k}^{\top}(\bm{V}_{i,k}+\bm{H}_{i,k}\bm{\Delta}_{k}\bm{H}_{i,k}^{\top})^{-1}\bm{H}_{i,k}\bm{\Delta}_{k}\bm{\Phi}^{\top}\right]$
with the initial condition $\bm{\Delta}_{0}=\bm{\Sigma}_{0}$.
Proof.
See [8, Thm. 3]. ∎
Remark 3.
Although rigorous, this bound is recursive and computationally expensive, providing certification but limited design insight. In particular, the absence of a closed-form steady-state solution prevents us from addressing questions such as how performance improves with additional sensors. This limitation motivates the continuous-time analysis in the next section, where explicit closed-form performance limits can be derived.
V-E2 Continuous-time Upper Bound for STGPKF
To overcome the limitations of the discrete-time recursion, we now turn to a continuous-time formulation. Following the approach in [11], we express the continuous-time measurement model corresponding to the discrete formulation in Eq. (LABEL:eq:measurement_model) as
$\bm{y}(t)=\bm{H}(\bm{r}(t))\bm{s}(t)+\bm{v}(t),\,\bm{v}(t)\sim\mathcal{N}\big(\bm{0},\bm{V}(\bm{r}(t))\big),$ (59)
where
$\bm{H}(t)$ $=\bm{K}_{rg}(\bm{r}(t))\bm{K}_{gg}^{-1}\sqrt{\bm{K}_{gg}}\bm{C},$
$\bm{V}(t)$ $=\Delta t\bm{V}_{k}.$
Combined with the continuous-time process model in Eq. (37), the covariance evolves according to the Kalman–Bucy Riccati differential equation in Eq. (5). The results developed in Theorems 1 and 2 can then be directly applied to characterize the expected estimation covariance and its steady-state behavior in the continuous-time setting.
Remark 4.
In simulation, however, it is important to note that the continuous-time filter cannot be directly realized. Instead, it is numerically approximated by discretizing time with step size $\Delta t$ and using the piecewise-constant information matrices constructed from $\bm{G}(t_{k})$ to represent the continuous matrix $\bm{G}(t)$. Only in the limit of sufficiently small sampling interval $\Delta t$ does the continuous-time filter coincide with the discrete-time filter. Nevertheless, the continuous-time formulation remains highly valuable: it provides an explicit, grid-independent characterization that admits closed-form solutions and exposes the fundamental sensing quantity that governs performance. We will show its advantages in the next section. Moreover, numerical results confirm that its relative deviation from the discrete-time bound is small, validating its use as an analytically tractable surrogate.
Figure 3: Comparison of continuous-time $\bm{\Delta}(t)$ and discrete-time $\bm{\Delta}_{k}$ covariance upper bounds with respect to the temporal resolution $\Delta t$.
To evaluate the effectiveness of our continuous-time bound, we compute the discrete-time expected covariance $\mathbb{E}[\bm{\Sigma}_{k}]$ by running the STGPKF over $30$ Monte Carlo trials with sensor locations randomized at each step, and compare it against (i) the discrete-time upper bound $\bm{\Delta}_{k}$ adapted from [8] and (ii) our analytically derived continuous-time upper bound $\bm{\Delta}(t)$. For visualization, we apply the linear operator $\mathcal{L}(\cdot)$ that maps a covariance matrix to the mean of its diagonal entries, and we report both the absolute error $\mathcal{L}(\bm{\Delta}_{k})-\mathcal{L}(\bm{\Delta}(t))$ and the relative error $\frac{\mathcal{L}(\bm{\Delta}_{k})-\mathcal{L}(\bm{\Delta}(t))}{\mathcal{L}(\bm{\Delta}_{k})}$ over time.
Figure 3 illustrates that as the sampling interval $\Delta t$ decreases, both the discrete-time and the continuous-time bounds converge to similarly low covariance levels. Moreover, the error between the two diminishes as the temporal resolution increases (i.e., smaller $\Delta t$). This trend confirms that the continuous-time formulation consistently approximates the discrete-time upper bound in the limit of small $\Delta t$, thus validating the continuous-time framework as a tractable surrogate for the discrete-time result.
V-E3 Continuous-time Lower Bound of Averaged Expected Clarity
Similar to the expected covariance, under stochastic sensing we further define the spatially averaged expected clarity as
$\bar{q}_{\mathbb{E}[\Pi]}(t):=\frac{1}{N_{g}}\sum_{i=1}^{N_{g}}\frac{1}{1+\mathbb{E}[\bm{\Pi}_{ii}(t)]}.$ (60)
where $\mathbb{E}[\bm{\Pi}(t)]=\mathcal{T}(\mathbb{E}[\bm{\Sigma}(t)])$ denotes expected covariance of field state.
Building on the covariance bound established earlier, we now seek a corresponding relation for the clarity in Eq. (60).
Theorem 3.
[Lower Bound on Averaged Expected Clarity] Suppose $\bm{\Delta}(t)\succeq\mathbb{E}[\bm{\Sigma}(t)]$ for all $t\geq t_{0}$. Then, the averaged expected clarity admits the lower bound
$\bar{q}_{\Delta^{\Pi}(t)}:=\frac{1}{1+\frac{1}{N_{g}}\operatorname{tr}(\bm{\Delta}^{\Pi}(t))},$ (61)
such that, for every $t\geq t_{0}$,
$\bar{q}_{\mathbb{E}[\Pi]}(t)\;\geq\;\bar{q}_{\Delta^{\Pi}(t)}.$ (62)
where $\bm{\Delta}^{\Pi}(t)=\mathcal{T}\big(\bm{\Delta}(t)\big)$ is the upper bound of the expected covariance of field state $\bm{f}$.
Proof.
Follows from Corollary 5, if $\bm{\Delta}(t)\succeq\mathbb{E}[\bm{\Sigma}(t)]$ then $L(\mathbb{E}[\bm{\Sigma}(t)])\preceq L(\bm{\Delta}(t))$. Since $\mathcal{T}(X)$ defined in Eq. (48) is linear and positive-semidefinite preserving, we obtain $\mathbb{E}[\bm{\Pi}(t)]=\mathcal{T}(\mathbb{E}[\bm{\Sigma}(t)])\preceq\mathcal{T}(\bm{\Delta}(t))=\bm{\Delta}^{\Pi}(t)$.
Let $f_{q}(x)=\frac{1}{1+x}$. Since $f_{q}$ is convex and strictly decreasing for $x\geq 0$, it follows that
$\frac{1}{N_{g}}\sum_{i=1}^{N_{g}}f_{q}\big(\mathbb{E}[\bm{\Pi}_{ii}(t)]\big)$ $\geq\frac{1}{N_{g}}\sum_{i=1}^{N_{g}}f_{q}(\bm{\Delta}^{\Pi}_{ii}(t))$ (63a)
$\geq f_{q}(\frac{1}{N_{g}}\sum_{i=1}^{N_{g}}\bm{\Delta}^{\Pi}_{ii}(t))$ (63b)
$=f_{q}(\frac{1}{N_{g}}\operatorname{tr}(\bm{\Delta}^{\Pi}(t)))$ (63c)
Eq. (63a) follows from the monotonicity of $f_{q}$ and the relation $\bm{\Delta}^{\Pi}(t)\succeq\mathbb{E}[\bm{\Pi}(t)]$, which implies $\bm{\Delta}^{\Pi}_{ii}(t)\geq\mathbb{E}[\bm{\Pi}_{ii}(t)]$ for all $i\in\{1,\dots,N_{g}\}$. Eq. (63b) follows from Jensen’s inequality, since $f_{q}$ is convex. Eq. (63c) uses the definition of the matrix trace, $\operatorname{tr}(\bm{\Delta}^{\Pi}(t))=\sum_{i=1}^{N_{g}}\bm{\Delta}^{\Pi}_{ii}(t)$. ∎
At steady state, we can get a closed-form expression for the lower bound:
$\bar{q}_{\Delta^{\Pi}_{\infty}}=\frac{1}{1+\frac{1}{N_{g}}\operatorname{tr}(\bm{\Delta}^{\Pi}_{\infty})}.$ (64)
which can be directly computed from the steady-state upper-bound covariance $\bm{\Delta}^{\Pi}_{\infty}$. As $\bm{\Delta}^{\Pi}_{\infty}$ depends on the eigenvalues of the averaged information matrix $\bar{\bm{G}}$ by Theorem 2, we next examine the structure of $\bar{\bm{G}}$.
V-F Analytical Characterization and Fundamental Limits
Building on the steady-state clarity bound derived in the previous section, this section addresses the second objective: to characterize the fundamental limits of estimation performance under stochastic sensing.
First, we derive a closed-form expression for the averaged information matrix $\bar{\bm{G}}$. Then, motivated by prior work [19], which investigates how system size affects estimation performance, we analyze the structure of the averaged information matrix $\bar{\bm{G}}$ to identify the factors fundamentally governing estimation accuracy. We aim to show that varying the number of grid points within the same compact domain does not fundamentally affect estimation performance; rather, it is dominated by intrinsic sensing parameters. As the grid resolution refines ($\delta\to 0$), the steady-state lower bound of the averaged expected clarity $\bar{q}_{\Delta^{\Pi}_{\infty}}$ becomes independent of discretization and, for a given temporal kernel, depends mainly on the number of agents, measurement noise, and sampling frequency.
V-F1 Closed-form of Averaged Information Matrix
Lemma 4.
If sensors take measurements at the grid points (i.e., sensing locations coincide with the discretization grid), then
$\bm{K}_{rr}-\bm{K}_{rg}\bm{K}_{gg}^{-1}\bm{K}_{gr}=0$ (65)
Proof.
See Appendix A-D. ∎
Lemma 5.
Let $P_{g}$ be as defined in Eq. (36) and suppose $N_{r}$ agents independently choose locations randomly from $P_{g}$. Then
$\mathbb{E}[\bm{K}_{gr}\bm{K}_{rg}]=\frac{N_{r}}{N_{g}}\bm{K}_{gg}^{2}$ (66)
Proof.
See Appendix A-E. ∎
Combining Lemma 4 and Lemma 5, we obtain the following expression for the averaged gain matrix $\bar{\bm{G}}$:
Lemma 6.
Define the sensing parameter
$\theta:=\frac{N_{r}}{\sigma_{m}^{2}\,\Delta t},$ (67)
where $N_{r}$ denotes the number of sensing agents, $\sigma_{m}^{2}$ is the measurement noise variance, and $\Delta t$ is the measurement interval. Under the same assumptions as in Lemma 4 and Lemma 5, for $N_{g}$ grid points and $N_{r}$ agents, the averaged information matrix has the following expression:
$\bar{\bm{G}}$ $=\mathbb{E}[\bm{H}\bm{V}^{-1}\bm{H}]$ (68)
$=\left(\frac{N_{r}}{\sigma_{m}^{2}\Delta t}\right)\bm{C}^{\top}\frac{\bm{K}_{gg}}{N_{g}}\bm{C}$
$=\theta\bm{C}^{\top}\frac{\bm{K}_{gg}}{N_{g}}\bm{C}$
Proof.
See Appendix A-F. ∎
It is observed that the averaged information matrix $\bar{\bm{G}}$ naturally separates the influence of the sensing configuration from that of the system dynamics. The scalar factor $\theta=\frac{N_{r}}{(\sigma_{m}^{2}\Delta t)}$ quantifies the sensing strength, jointly determined by the number of sensors, their measurement noise level, and the measurement rate. In contrast, the matrix term $\bm{C}^{\top}\frac{\bm{K}_{gg}}{N_{g}}\bm{C}$ encapsulates the spatial correlation structure of the environment. Moreover, since $\bar{\bm{G}}$ is mainly determined by the spatial kernel matrix $\bm{K}_{gg}$, its spectral properties govern how sensing resources affect estimation clarity. This observation motivates a spectral analysis of $\bm{K}_{gg}$ in the next subsection.
V-F2 Fundamental Performance Limits
In this section, we investigate how the system size, represented by the number of grid points to be estimated, affects the averaged estimation performance. Specifically, we will describe how the discrete spatial kernel matrix $\bm{K}_{gg}$ behaves as the grid resolution becomes arbitrarily fine.
We first require spectral properties of the kernel integral operator, stated in the following lemma.
Lemma 7 (Mercer’s theorem).
Consider a continuous, symmetric, positive semi-definite kernel $k_{S}$ defined in Eq. (33) on a compact domain $D$. Assume the operator $T$ takes a function $f(x)$ as its argument and outputs a new function as:
$Tf(x):=\int_{D}k(x,y)f(y)\,dy,$ (69)
The operator $T$ is called the Hilbert–Schmidt integral operator. Then, there is a set of orthonormal bases $\{\psi_{i}(\cdot)\}_{i=1}^{\infty}$ of $L_{2}(D)$ consisting of eigenfunctions of $T$ such that the corresponding sequence of eigenvalues $\{\nu_{i}\}_{i=1}^{\infty}$ are non-negative and
$k(x,y)=\sum_{i=1}^{\infty}\nu_{i}\psi_{i}(x)\psi_{i}(y),$ (70)
where the convergence is absolute and uniform. Moreover, $T$ is trace class with
$\sum_{i=1}^{\infty}\nu_{i}=\int_{D}k_{S}(x,x)\,dx<\infty.$ (71)
Proof.
See [23, Thm. 2] and [24, Cor. 5.4,Thm 5.2]. ∎
Notice that the spatial kernel matrix $\bm{K}_{gg}$ defined in Eq. (38) is a Nyström approximation of the continuous kernel operator $T$. This approximation is constructed by sampling the continuous kernel function $k_{S}$ on a uniform spatial grid with spacing $\delta$. The following lemma shows that as $\delta\to 0$, the eigenvalues of discrete operator $T_{\delta}$ converge to those of $T$.
Lemma 8.
Let $D\subset\mathbb{R}^{d}$ be a compact domain with finite Lebesgue measure $|D|>0$. For a uniform grid spacing $\delta>0$, define the Nyström approximation operator:
$(T_{\delta}f)(p):=\frac{|D|}{N_{g}}\sum_{j=1}^{N_{g}}k_{S}(p,p_{j})f(p_{j}),N_{g}=\frac{|D|}{\delta^{d}}$ (72)
where $\{p_{j}\}_{j}^{N_{g}}\subset D$ are grid points and $k_{S}$ is a continuous, symmetric, positive-definite kernel defined in Eq. (33). Let $\{\nu_{i,\delta}\}_{i=1}^{N_{g}}$ and $\{\nu_{i}\}_{i=1}^{\infty}$ denote eigenvalues of $T_{\delta}$ and of continuous operator $T$ defined in Lemma 7, respectively. Then, as $\delta\to 0$ ($N_{g}\to\infty$),
$\nu_{i,\delta}\to\nu_{i},\,\forall i\geq 1.$ (73)
Proof.
See [25, Thm. 4.2]. ∎
We can now quantify the asymptotic behavior of $\bar{q}_{\Delta^{\Pi}_{\infty}}$ as the grid becomes dense, which constitutes one of our main results.
Theorem 4.
Consider the Matérn-$\tfrac{1}{2}$ kernel ($n_{k}=1$) with $\bm{C}=C_{0}\bm{I}$ and eigenvalues $\{\nu_{i}\}_{i=1}^{\infty}$ of the associated continuous kernel operator in Eq. (33) on a given spatial domain. Then, as the grid spacing $\delta\to 0$, the steady-state lower bound of the averaged expected clarity defined in Eq. (64) converges to a finite limit, i.e.,
$\lim_{\delta\to 0}\bar{q}_{\Delta^{\Pi}_{\infty}}=\frac{1}{1+C_{0}^{\,2}\sum_{i=1}^{\infty}\nu_{i}\gamma\bigl(\theta C_{0}^{2}\nu_{i}\bigr)}$ (74)
which depends solely on the sensing parameter
$\theta=\frac{N_{r}}{\sigma^{2}\Delta t}.$ (75)
Proof.
Let the normalized spatial kernel matrix be
$\frac{\bm{K}_{gg}}{N_{g}}=\bm{U}\,\mathrm{diag}(\nu_{1,\delta},\ldots,\nu_{N_{g},\delta})\,\bm{U}^{\top}=\bm{U}\,\frac{\bm{\Lambda}_{K}}{N_{g}}\,\bm{U}^{\top}$ (76)
where $\bm{U}$ is orthogonal and $\nu_{i,\delta}$ are the eigenvalues of $\frac{\bm{K}_{gg}}{N_{g}}$. From Eq. (68),
$\bar{\bm{G}}=\theta C_{0}^{2}\bm{U}\frac{\Lambda_{K}}{N_{g}}\bm{U}^{T}=\bm{U}\mathrm{diag}\left(\theta C_{0}^{2}\nu_{1,\delta},\ldots,\theta C_{0}^{2}\nu_{N_{g},\delta}\right)\bm{U}^{\top}$ (77)
Define $\bm{\Lambda}_{\Delta}=\mathrm{diag}\big(\gamma(\theta C_{0}^{2}\nu_{1,\delta}),\ldots,\gamma(\theta C_{0}^{2}\nu_{N_{g},\delta})\big)$. According to Theorem 2,
$\bm{\Delta}_{\infty}=\bm{U}\mathrm{diag}\big(\gamma(\theta C_{0}^{2}\nu_{1,\delta}),\ldots,\gamma(\theta C_{0}^{2}\nu_{N_{g},\delta})\big)\bm{U}^{\top}.$ (78)
The corresponding normalized upper bound of steady-state covariance of field state is
$\frac{\bm{\Delta}^{\Pi}_{\infty}}{N_{g}}$ $=\frac{1}{N_{g}}C_{0}^{2}\bm{K}_{gg}^{1/2}\bm{\Delta}_{\infty}\bm{K}_{gg}^{1/2}$
$=\frac{1}{N_{g}}C_{0}^{2}(\bm{U}\bm{\Lambda}_{K}^{1/2}\bm{U}^{\top})(\bm{U}\bm{\Lambda}_{\Delta}\bm{U}^{\top})(\bm{U}\bm{\Lambda}_{K}^{1/2}\bm{U}^{\top})$
$=\frac{1}{N_{g}}C_{0}^{2}\bm{U}(\bm{\Lambda}_{K}^{1/2}\bm{\Lambda}_{\Delta}\bm{\Lambda}_{K}^{1/2})\bm{U}^{\top},$
$=C_{0}^{2}\bm{U}(\frac{\bm{\Lambda}_{K}}{N_{g}}\bm{\Lambda}_{\Delta})\bm{U}^{\top}.$
Therefore, the eigenvalues of $\frac{\bm{\Delta}^{\Pi}_{\infty}}{N_{g}}$ are $\{\nu_{i,\delta}\gamma(\theta C_{0}^{2}\nu_{i,\delta})\}_{i}^{N_{g}}$. Hence,
$\frac{1}{N_{g}}\operatorname{tr}(\bm{\Delta}^{\Pi}_{\infty})=\sum_{i}^{N_{g}}\lambda_{i}(\frac{\bm{\Delta}^{\Pi}_{\infty}}{N_{g}})=\sum_{i}^{N_{g}}C_{0}^{2}\nu_{i,\delta}\gamma(\theta C_{0}^{2}\nu_{i,\delta})$
By Mercer’s theorem in Lemma 7 and its Nyström approximation in Lemma 8, $\{\nu_{i,\delta}\}_{i=1}^{N_{g}}$ are approximations of the eigenvalues $\{\nu_{i}\}_{i=1}^{\infty}$ of the continuous kernel operator, satisfying $\nu_{i,\delta}\to\nu_{i}$ as $\delta\to 0$ and $\sum_{i=1}^{\infty}\nu_{i}<\infty$. From Theorem 2, $\gamma(\cdot)$ is continuous and bounded, i.e.,
$0\leq\gamma(\lambda)\leq\gamma(0)=:C_{\gamma}<\infty,$ (79)
we have
$0\leq\nu_{i}\gamma(\theta C_{0}^{2}\nu_{i})\leq C_{\gamma}\nu_{i}.$ (80)
Therefore,
$\lim_{\delta\to 0}\frac{1}{N_{g}}\,\operatorname{tr}\bigl(\bm{\Delta}^{\bm{\Pi}}_{\infty}\bigr)$ $=C_{0}^{\,2}\sum_{i=1}^{\infty}\nu_{i}\gamma\bigl(\theta C_{0}^{2}\nu_{i}\bigr)$
$\leq C_{0}^{\,2}\sum_{i=1}^{\infty}\nu_{i}C_{\gamma}$
converges to a finite constant.
Since $f_{q}=\frac{1}{1+x}$ is bounded and continuous on $0,\infty)$, the steady–state lower bound
$\lim_{\delta\to 0}\bar{q}_{\Delta^{\Pi}_{\infty}}$ $=\frac{1}{1+\lim_{\delta\to 0}\frac{1}{N_{g}}\operatorname{tr}(\bm{\Delta}^{\Pi}_{\infty})}$
$=\frac{1}{1+C_{0}^{\,2}\sum_{i=1}^{\infty}\nu_{i}\gamma\bigl(\theta C_{0}^{2}\nu_{i}\bigr)}$
converges as $\delta\to 0$. When the environment (kernel parameters and domain) is fixed, the steady-state lower bound converges to a constant that depends solely on the sensing parameter $\theta$, independent of the grid resolution. ∎
Remark 5.
Theorem 4 shows that $\bar{q}_{\Delta^{\Pi}_{\infty}}$ converges as $\delta\to 0$, implying that beyond a certain system size, adding more measurement points no longer improves estimation performance. When the temporal kernel parameters $(\bm{A}_{0},\bm{B}_{0},\bm{C}_{0})$ are fixed, this limit depends only on the sensing parameter $\theta$ which combines the number of agents $N_{r}$, the measurement noise variance $\sigma_{m}^{2}$, and the sampling interval $\Delta t$. We refer to $\theta$ as the fundamental quantity governing sensing performance.
Figure 4: Steady-state lower bound of averaged expected clarity $\bar{q}_{\Delta^{\Pi}_{\infty}}$ versus spatial grid size $N_{g}$. Each curve corresponds to a different $\theta(N_{r})$, representing a different number of sensors.
Figure 4 empirically validates Theorem 4. The curves are parameterized by $\theta$, showing that the steady-state of lower bound of averaged expected clarity $\bar{q}_{\Delta^{\Pi}_{\infty}}$ converges to distinct steady-state values determined by $\theta$, which characterizes the fundamental performance limit. This confirms that estimation performance is dominated by intrinsic sensing parameters rather than spatial discretization.
V-F3 Number of Sensors
Two main consequences follow. First, the limiting performance is intrinsic to the kernel and system dynamics rather than an artifact of grid resolution. Second, the dependence on a single parameter $\theta$ provides a clear design principle for a given environment setting: performance can be improved equivalently by adding agents, reducing measurement noise, or increasing sampling frequency. Moreover, the minimum sensor number required to meet a desired performance threshold can be computed directly based on Theorem 4.
We now revisit the design question: Given environment parameters, how many sensors are needed to guarantee a desired uncertainty threshold?
The inequality in Eq. (62) implies that the averaged expected clarity $\bar{q}_{\mathbb{E}[\Pi]}(t)$ can be certified by evaluating it on the upper bound covariance $\bm{\Delta}(t)$, which gives lower bound of averaged expected clarity $\bar{q}_{\Delta}(t)$. Enforcing clarity constraints for all $t$ is intractable. However, Theorem 4 ensures that $\bar{q}_{\Delta}(t)$ admits a well-defined steady-state limit as the grid is refined. Using this steady-state clarity $\bar{q}_{\Delta^{\Pi}_{\infty}}$ as a surrogate is both tractable and sufficient, as it captures the long-term behavior of the KF and guarantees the desired level of expected clarity. Consequently, the agent-number design problem can be formulated as the following optimization:
$\min\;N_{r}\quad\text{s.t.}\;\bar{q}_{\Delta^{\Pi}_{\infty}}\geq q_{\text{target}}.$ (81)
where $q_{\text{target}}$ denotes the user-defined clarity level.
TABLE III: $\bar{q}_{\mathbb{E}[\Pi]}$ and $\bar{q}_{\Delta^{\Pi}_{\infty}}$ versus number of sensors.
$q_{\text{target}}$ $N_{r}$ $\bar{q}_{\mathbb{E}[\Pi]}$ $\bar{q}_{\Delta^{\Pi}_{\infty}}$
0.50 1 0.511 0.528
0.60 3 0.621 0.632
0.70 7 0.704 0.712
0.80 21 0.798 0.803
0.90 112 0.897 0.900
Figure 5: $\bar{q}_{\mathbb{E}[\Pi]}$ and $\bar{q}_{\Delta^{\Pi}_{\infty}}$ versus the number of sensors.
Results show that achievable clarity increases monotonically with $N_{r}$, though with diminishing returns as the number of sensors grows. Crucially, across all tested configurations, $\bar{q}_{\mathbb{E}[\Pi]}$ consistently satisfies the target threshold $q_{\text{target}}$, confirming that the design strategy in Eq. (81) effectively identifies the minimal number of agents needed to guarantee a desired clarity level. More importantly, there is a practical guideline: given a target clarity level, one can directly read from Figure 5 and Table III the minimum number of agents required to achieve it. The close agreement between $\bar{q}_{\mathbb{E}[\Sigma]}$ and $\bar{q}_{\Delta^{\Pi}_{\infty}}$ further validates that $\bar{q}_{\Delta^{\Pi}_{\infty}}$ provides a reliable surrogate for the sensor-number design problem.
With the number of sensors fixed, we can also examine how sensing capabilities affect performance. Figure 6 shows that the covariance decreases smoothly from the lower-left to the upper-right: smaller $\theta$ corresponds to harsher sensing conditions, i.e., higher noise or slower sampling. Here, $\theta:=\tfrac{N_{r}}{\sigma_{m}^{2}\Delta t}$ quantifies sensing intensity. Moving along any dashed constant sensing intensity (constant $\theta$) curve leaves the mean clarity bound nearly unchanged. Thus, selecting a target clarity determines a required $\theta$, and any $(\sigma_{m}^{2},\Delta t)$ pair on the constant $\theta$ curve (for the fixed $N_{r}$) guarantees the same performance. This characterization enables explicit trade-offs between sensor noise and measurement rate prior to deployment.
Figure 6: Steady-state lower bound of averaged expected clarity $\bar{q}_{\Delta^{\Pi}_{\infty}}$ as a function of sensing noise and measurement rate given fixed $N_{r}=1$.
VI Conclusions
This paper developed a continuous-time framework for quantifying the fundamental performance limits of multi-agent spatiotemporal estimation under stochastic sensing. We derived the first continuous-time upper bound on the expected Kalman–Bucy covariance under randomized sensing, extending discrete-time results to a broader analytical setting. Applied to spatiotemporal field estimation, the framework yields a grid-independent performance law governed by a parameter combining agent number, measurement noise, and measurement rate. Simulation results show that the continuous-time bound accurately predicts the empirical estimation performance and more importantly, it provides a reliable tool for pre-deployment sensor number design without costly simulations. Future research will consider incorporating the dynamics and motion constraints of mobile sensors into the estimation framework and address non-uniform location sampling arising from heterogeneous sensor distributions.
Appendix A Proofs
A-A Proof of Lemma 1
Proof.
Let
$\bm{\Sigma}_{\lambda}=\lambda\bm{\Sigma}_{1}+(1-\lambda)\bm{\Sigma}_{2}$
with $\lambda\in[0,1]$.
The affine part is linear, so
$\bm{A}\bm{\Sigma}_{\lambda}+\bm{\Sigma}_{\lambda}\bm{A}^{\top}+\bm{Q}_{c}$
$=\lambda(\bm{A}\bm{\Sigma}_{1}+\bm{\Sigma}_{1}\bm{A}^{\top}+\bm{Q}_{c})+(1-\lambda)(\bm{A}\bm{\Sigma}_{2}+\bm{\Sigma}_{2}\bm{A}^{\top}+\bm{Q}_{c})$
For the quadratic term,
$\bm{\Sigma}_{\lambda}\bm{G}\bm{\Sigma}_{\lambda}$
$=(\lambda\bm{\Sigma}_{1}+(1-\lambda)\bm{\Sigma}_{2})\bm{G}(\lambda\bm{\Sigma}_{1}+(1-\lambda)\bm{\Sigma}_{2})$
$=\lambda^{2}\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{1}+(1-\lambda)^{2}\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{2}$
$\quad+\lambda(1-\lambda)(\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{2}+\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{1})$
$=\lambda\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{1}+(1-\lambda)\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{2}$
$\quad-\lambda(1-\lambda)(\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{1}+\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{2}-(\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{2}+\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{1}))$
$=\lambda\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{1}+(1-\lambda)\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{2}$
$\quad-\lambda(1-\lambda)(\bm{\Sigma}_{1}-\bm{\Sigma}_{2})\bm{G}(\bm{\Sigma}_{1}-\bm{\Sigma}_{2})$
Since $\bm{G}\succeq 0$, we have $(\bm{\Sigma}_{1}-\bm{\Sigma}_{2})\bm{G}(\bm{\Sigma}_{1}-\bm{\Sigma}_{2})\geq 0$.
$\bm{\Sigma}_{\lambda}\bm{G}\bm{\Sigma}_{\lambda}$
$\leq\lambda\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{1}+(1-\lambda)\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{2}$
Hence,
$\mathcal{R}_{\bm{G}}(\lambda\bm{\Sigma}_{1}+(1-\lambda)\bm{\Sigma}_{2})$
$=\bm{A}\bm{\Sigma}_{\lambda}+\bm{\Sigma}_{\lambda}\bm{A}^{\top}+\bm{Q}_{c}-\bm{\Sigma}_{\lambda}\bm{G}\bm{\Sigma}_{\lambda}$
$\geq\lambda(\bm{A}\bm{\Sigma}_{1}+\bm{\Sigma}_{1}\bm{A}^{\top}+\bm{Q}_{c})+\lambda\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{1}$
$\quad+(1-\lambda)(\bm{A}\bm{\Sigma}_{2}+\bm{\Sigma}_{2}A^{\top}+\bm{Q}_{c})+(1-\lambda)\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{2}$
$=\lambda\mathcal{R}_{\bm{G}}(\bm{\Sigma}_{1})+(1-\lambda)\mathcal{R}_{\bm{G}}(\bm{\Sigma}_{2})$
A-B Proof of Lemma 2
Proof.
The proof is inspired by [26, Theorem 2.1].
$\mathcal{R}_{\bm{G}}(\bm{\Sigma}_{1})-\mathcal{R}_{\bm{G}}(\bm{\Sigma}_{2})$
$=\bm{A}\bm{\Sigma}_{1}+\bm{\Sigma}_{1}\bm{A}^{\top}-\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{1}-\bm{A}\bm{\Sigma}_{2}-\bm{\Sigma}_{2}\bm{A}^{\top}+\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{2}$
$=\bm{A}(\bm{\Sigma}_{1}-\bm{\Sigma}_{2})+(\bm{\Sigma}_{1}-\bm{\Sigma}_{2})\bm{A}^{\top}-\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{1}+\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{2}$
$=AK+\bm{K}\bm{A}^{\top}-\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{1}+\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{2}$
$=AK+\bm{K}\bm{A}^{\top}-\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{1}+(\bm{\Sigma}_{1}-\bm{K})\bm{G}(\bm{\Sigma}_{1}-\bm{K})$
$=AK+\bm{K}\bm{A}^{\top}-\bm{\Sigma}_{1}\bm{G}\bm{K}-\bm{K}\bm{G}\bm{\Sigma}_{1}+\bm{K}\bm{G}\bm{K}$
$=(\bm{A}-\bm{\Sigma}_{1}\bm{G}+\frac{1}{2}\bm{K}\bm{G})\bm{K}+\bm{K}(\bm{A}^{\top}-\bm{G}\bm{\Sigma}_{1}+\frac{1}{2}\bm{G}\bm{K})$
$=(\bm{A}-\frac{1}{2}(\bm{\Sigma}_{1}+\bm{\Sigma}_{2})\bm{G})\bm{K}+\bm{K}(\bm{A}^{\top}-\frac{1}{2}\bm{G}(\bm{\Sigma}_{1}+\bm{\Sigma}_{2}))$
$=\tilde{\bm{A}}\bm{K}+\bm{K}\tilde{\bm{A}}^{\top}$.
A-C Proof of Lemma 3
Proof.
Let $\bm{\Psi}(t,t_{0})\in\mathbb{R}^{n\times n}$ be the state-transition matrix associated with $\tilde{\bm{A}}^{\top}(t)$, i.e.,
$\dot{\bm{\Psi}}(t,t_{0})=\tilde{\bm{A}}^{\top}(t)\bm{\Psi}(t,t_{0}),\quad\bm{\Psi}(t_{0},t_{0})=\bm{I}.$
Define
$\bm{M}(t):=\bm{\Psi}(t,t_{0})^{-T}\bm{K}(t)\bm{\Psi}(t,t_{0})^{-1}.$
Using $\frac{d}{dt}\bm{\Psi}^{-1}=-\bm{\Psi}^{-1}\tilde{\bm{A}}^{\top}$ and $\frac{d}{dt}\bm{\Psi}^{-\top}=-\tilde{\bm{A}}\bm{\Psi}^{-\top}$, the product rule gives
$\dot{\bm{M}}$ $=\frac{d}{dt}\left(\bm{\Psi}^{-T}\right)\bm{K}\bm{\Psi}^{-1}+\bm{\Psi}^{-T}\dot{\bm{K}}\bm{\Psi}^{-1}+\bm{\Psi}^{-T}\bm{K}\frac{d}{dt}\left(\bm{\Psi}^{-1}\right)$
$=\bm{\Psi}^{-T}(\dot{\bm{K}}-\tilde{\bm{A}}(t)\bm{K}-\bm{K}\tilde{\bm{A}}(t)^{\top})\bm{\Psi}^{-1}\succeq 0$
where we omitted $(t,t_{0})$ for brevity. Hence, $\bm{M}(\cdot)$ is monotone nondecreasing in the Loewner order and
$\bm{M}(t)\succeq\bm{M}(t_{0})=\bm{K}(t_{0})\succeq 0,\quad\forall t\geq t_{0}.$
Finally, by congruence, we have
$\bm{K}(t)=\bm{\Psi}(t,t_{0})^{\top}\bm{M}(t)\bm{\Psi}(t,t_{0})\succeq 0,\quad\forall t\geq t_{0}.$
A-D Proof of Lemma 4
Proof.
Define the selection matrix $\bm{S}\in\mathbb{R}^{N_{g}\times N_{r}}$ is a column selector matrix, indicating which grid point the agent is at. We have
$\bm{K}_{rr}=\bm{S}^{\top}\bm{K}_{gg}\bm{S},\,\bm{K}_{rg}=\bm{S}^{\top}\bm{K}_{gg},\,\bm{K}_{gr}=\bm{K}_{gg}\bm{S}$
Therefore,
$\bm{K}_{rg}\bm{K}^{-1}_{gg}\bm{K}_{gr}$
$=(\bm{S}^{\top}\bm{K}_{gg})\bm{K}^{-1}_{gg}(\bm{K}_{gg}\bm{S})$
$=\bm{S}^{\top}\bm{I}\bm{K}_{gg}\bm{S}$
$=\bm{K}_{rr}$
Thus,
$\bm{K}_{rr}-\bm{K}_{rg}\bm{K}_{gg}^{-1}\bm{K}_{gr}=0$
A-E Proof of Lemma 5
Proof.
Since the agents are at one of the grid points, we can write
$\bm{K}_{gr}=\bm{K}_{gg}\bm{S}$
where $\bm{S}\in\mathbb{R}^{N_{g}\times N_{r}}$ is a column selector matrix, indicating which gridpoint the agent is at. Each column of $\bm{S}$ is an element of standard basis set. Therefore,
$\mathbb{E}[\bm{K}_{gr}\bm{K}_{rg}]$ $=\mathbb{E}[\bm{K}_{gg}\bm{S}\bm{S}^{\top}\bm{K}_{gg}]$
$=\bm{K}_{gg}\mathbb{E}[\bm{S}\bm{S}^{\top}]\bm{K}_{gg}$
$=\bm{K}_{gg}\left(\frac{N_{r}}{N_{g}}\bm{I}\right)\bm{K}_{gg}.$
A-F Proof of Lemma 6
Proof.
Since the agents take measurements at the grid locations, by Lemma 4, we have
$\bm{V}(\bm{r})$ $=(\Delta t)(\sigma_{m}^{2}\bm{I}_{N_{r}}+\bm{K}_{rr}-\bm{K}_{rg}\bm{K}_{gg}^{-1}\bm{K}_{gg})$
$=(\Delta t\sigma_{m}^{2})\bm{I}_{N_{r}}.$
Moreover, by Lemma 5, we have
$\mathbb{E}[\bm{H}^{\top}\bm{V}^{-1}\bm{H}]$
$=\frac{1}{\Delta t\sigma_{m}^{2}}\mathbb{E}[\bm{H}^{\top}\bm{H}]$
$=\frac{1}{\Delta t\sigma_{m}^{2}}\mathbb{E}[\bm{C}^{\top}\bm{K}_{gg}^{-1/2}\bm{K}_{gr}\bm{K}_{rg}\bm{K}_{gg}^{-1/2}\bm{C}]$
$=\frac{1}{\Delta t\sigma_{m}^{2}}\bm{C}^{\top}\bm{K}_{gg}^{-1/2}\mathbb{E}[\bm{K}_{gr}\bm{K}_{rg}]\bm{K}_{gg}^{-1/2}\bm{C}$
$=\frac{1}{\Delta t\sigma_{m}^{2}}\bm{C}^{\top}\frac{N_{r}}{N_{g}}\bm{K}_{gg}\bm{C}$
$=\frac{N_{r}}{\Delta t\sigma_{m}^{2}}\bm{C}^{\top}\frac{\bm{K}_{gg}}{N_{g}}\bm{C}$